OpenStudy (chupacabraj):

Find a vector function that represents the curve of intersection hyperboloid z=x^2-y^2 and the cylinder x^2+y^2=1

OpenStudy (chupacabraj):

I tried isolating the x in the second equation and then plugging that into the first one but it turned out pretty ugly and then I just didn't know where I was going :(

OpenStudy (3mar):

why you just did not sum them ?

OpenStudy (chupacabraj):

No, why should I?

OpenStudy (chupacabraj):

@mathmale

OpenStudy (otherworldly):

@jackthegreatest

OpenStudy (jackthegreatest):

i have no idea : (

OpenStudy (chupacabraj):

I really want to understand how to solve these types of problems.

OpenStudy (tkhunny):

It is not a plug-and-chug sort of thing. It is a matter of understanding what you are seeing. z=x^2-y^2 and the cylinder x^2+y^2=1 z = x^2 + y^2 - 2y^2 = 1 - 2y^2 Ever see one of these? z = 1 - 2y^2 These are cross-sections and they are parabolas. No reason the same sort of thing wouldn't work for eliminating the y^2 and seeing what other sorts of cross sections abound.

OpenStudy (chupacabraj):

what z = x^2 + y^2 - 2y^2 = 1 - 2y^2 ?

OpenStudy (chupacabraj):

Hey

OpenStudy (chupacabraj):

what did you add and subtract?

OpenStudy (tkhunny):

z = x^2 - y^2 Add and subtract y^2 z = x^2 + y^2 - y^2 - y^2 Simplify Conveniently z = x^2 + y^2 - 2y^2 Given that x^2 + y^2 = 1 z = 1 - 2y^2 Such manipulations are very helpful.

OpenStudy (tkhunny):

I added and subtracted because I already knew x^2 + y^2 = 1 and I needed that structure. So I built it.

OpenStudy (chupacabraj):

ohhhh makes sense

OpenStudy (chupacabraj):

Now, I can say if y= t, then z=1-2t^2 ?

OpenStudy (tkhunny):

I suppose. If that has some relevance to the present investigation.

OpenStudy (chupacabraj):

Well, I need a vector function so I have to "parametrize" right?

OpenStudy (3mar):

Thank you a lot, tkhunny. You made it so clear, for me at least.

OpenStudy (chupacabraj):

Whoah I'm confused, I checked the answer key and it says the vector function is cost i + sin t j + cos2t k how did it end up being trig?

OpenStudy (kevin):

I'm here just to saying "woah" xD

OpenStudy (tkhunny):

Cylinders have circular cross sections. You will get trigonometric functions.

OpenStudy (phi):

It makes sense to use cylindrical coordinates x= r cos t y = r sin t z = z in this case \[ x^2+y^2 = 1\\ r^2 \cos^2 t + r^2 \sin^2 t = 1\\ r=1\] and \[ z= x^2 -y^2 \\= r^2(\cos^2 t - \sin^2 t) \\ = r \cos 2t \\= \cos 2t \] we get \[ < \cos t , \sin t , \cos 2t >\]